# Understanding and Imitating Human-Robot Motion with Restricted Visual Fields

### v1

@@ Abstract
When working around humans, it is important to model their perception limitations in order to predict their behavior more accurately. In this work, we consider agents with a limited field of view, viewing range, and ability to miss objects within viewing range (e.g., transparency). By considering the observation model independently from the motion policy, we can better predict the agent’s behavior by considering these limitations and approximating them. We perform a user study where human operators navigate a cluttered scene while scanning the region for obstacles with a limited field of view and range. Using imitation learning, we show that a robot can adopt a human’s strategy for observing an environment with limitations on observation and navigate with minimal collision with dynamic and static obstacles. We also show that this learned model helps it successfully navigate a physical hardware vehicle in real time. Code available at https://github.com/labicon/HRMotion-RestrictedView.
@@ I INTRODUCTION
There has been extensive literature on learning from human demonstrations and imitating them. An agent’s policy can be approximated and replicated using methods from Learning from Demonstration [8, 9, 10]. By observing a demonstrator, the agent’s reward function can be inferred from their actions (Inverse Reinforcement Learning - IRL) [11]. Alternatively, one can directly learn and predict the actions of the agent using Behavior Cloning Strategies (BC) [12, 13, 14, 15] or hybrid approaches where a discriminator must distinguish between artificial actions from a generator and actual demonstration actions, and the generator is capable of sufficiently confusing the discriminator is a trained agent [16, 17]. When it comes to BC methods, Diffusion models have seen a recent rise in their application to learning from the demonstration as actions are diffused from noise conditioned on observations and states and accommodate multi-modality  [18, 19, 20, 21, 22]. While these models have proven to be successful in modeling human choices and intentions in a number of settings, we argue that to better model, predict, and imitate agents such as humans, we need to account for the perception capabilities and observation space of the agents as well. Leveraging these models, we argue that robots that can approximate the observation of another (human) agent, including the field of view, head/gaze/camera orientation, and the likelihood of observing an object, in addition to the motion policy, to better predict the actions of that agent since the agents’ motion policy will be conditioned on their observations.
@@
Our contributions are as follows: to better understand the relationship between an agent’s observation strategy and motion policy, we create a robotic agent with a limited field of view, view range, and likelihood of observing an obstacle and use the agent’s motion to infer these quantities. We consider three distinct scenarios, where each scenario involves one unknown: the parameters of the observation space, the probability of detecting obstacles within the agent’s field of view, or the agent’s motion policy, which depends on its observations. In all scenarios, we show that by accounting for the agent’s inherent observation limitations, we can more accurately model and predict its behavior. More specifically, we show that to infer the parameters of the observation space; one can use the cross-entropy [23] method to estimate the observation parameters accurately. We show that using Bayesian optimization [24], the probability of object detection within an agent’s field of view can be easily estimated. Finally, for inferring the policy of the agent, we show that the agent’s policy can be accurately imitated using diffusion models [19]. We illustrate this by training a diffusion policy on human demonstrations of navigating an environment under a restricted field of view. We demonstrate our ability to predict and replicate human behavior for both observation and motion in the presence of static and dynamic obstacles. Finally, we deploy the imitated policy to physical hardware, showcasing a robot navigating a cluttered environment and exhibiting human-like behavior by dynamically rotating their field of view to detect obstacles.
@@ II-A Problem Formulation
Consider an agent navigating in an environment with obstacles. The obstacles could be either stationary or moving. We use $x_{t}\in\mathbb{R}^{n}$ to denote the state of the agent at time $t\in\mathbb{N}\cup\{0\}$, where is =$\mathbb{N}=\{1,2,3,\ldots\}$ is the set of natural numbers. Note that here, $n$ represents the agent’s state space dimension. Let the control action of the agent at time $t$ be denoted by $u_{t}\in\mathbb{R}^{m}$ where $m$ is the dimension of the control space of the agent. We consider discrete deterministic dynamics for the agent denoted through a function $f:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$,
@@
We assume that we know the robot’s dynamics and that the agent has limited observation capabilities. The agent’s observation at time $t$ is denoted via $z_{t}\in\mathbb{R}^{o}$ where $o$ is the dimension of the observation space of the agent. We assume that the agent has an observation space that determines the area of the environment the agent is looking towards. Furthermore, we assume that the agent has an inherent probability of observing something present in its observation area. The combined observation space and probability of observation determine the agent’s observation at each time, given the agent’s current state. Let the observation space parameter that determines the agent’s observation area be denoted by $\theta\in\Theta$ where $\Theta$ is the space of observation space parameters. Also, let the probability of observing a particular observation be denoted by $p_{obs}\in[0,1]$. Therefore, given that the agent’s state is $x_{t}$, the probability of observing an observation $z_{t}$ is denoted via
@@
At each time $t$, based on the observation $z_{t}$, the agent takes a control action $u_{t}$. The probability of taking an action at each time $t$ by the agent is captured through the agent policy, which we denote via $\pi:\mathbb{R}^{m}\times\mathbb{R}^{o}\rightarrow[0,1]$,
@@
We use $\tau$ to denote a trajectory of the agent as $\tau=\{x_{0},u_{0},x_{1},u_{1},\ldots,x_{T-1},u_{T-1},x_{T}\}$, where $T$ is the length of the trajectory. In this work, we assume that we have access to a set of $N$ trajectories, $\mathcal{D}=\{\tau_{i}\}_{i=1}^{N}$ exhibited by the agent in an environment. We aim to estimate the agent’s observation space parameter $\theta$ and policy parameter $\theta_{\pi}$ based on the trajectory data to mimic human behaviors.
Running Example: To illustrate this idea, consider an agent with discrete dynamics (Eqn. (1)) navigating through an environment with obstacles to a goal (Figure 1). The agent visualizes the scene with a visual sensor parameterized by the range of vision ($r_{obs}$), and the field of view ($\theta_{obs}$). Obstacles in the field of view are not guaranteed to be seen (due to transparency and lack of visual features), which is captured by the detection probability $p_{obs}$. The motion of the agent is governed by the policy $\pi(\cdot;\theta_{\pi})$, which controls a) the motion of the robot in the environment, and b) the orientation of the visual sensor denoted by $\psi_{obs}$. With these abilities and observation limitations, the agent navigates the cluttered environment trying to minimize collision with obstacles while reaching the goal.
@@ II-C Estimating the Probability of Observation
$p_{obs}^{*}=\operatorname*{arg\,max}_{p_{obs}\in[0,1]}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\log(\int O(z_{t},x_{t};\theta,p_{obs})\pi(u_{t},z_{t};\theta_{\pi})).$ (6)
@@ III Numerical Simulations
In this section, we present numerical simulations to showcase the importance of considering agents’ observation and perception capabilities when either learning from them or modeling their behavior. In this section, we consider synthetic settings where we have access to the ground-truth parameters which enable a comprehensive comparison and evaluation of our method. Then, in the next section, we consider human data and perform hardware experiments that leverage actual human data. We consider a unicycle agent with discrete dynamics
@@
Figure 3: Schematic for the unicycle navigation scenario with partial observations. The red triangle represents the unicycle. The gray circles represent the environmental obstacles, and the green represents the goal. The orange cone represents the observation space of the agent, which is parameterized by the radius of the cone ($r_{obs}$) and the angle of the cone ($\theta_{obs}$). Each obstacle lights up red when the agent observes the obstacle.
@@
The Fréchet distances between actual and predicted trajectories for all scenarios are provided in Table II. As can be seen, for 3 of the four datasets, the estimated parameters are within about $5\%$ of the interval to the original trajectories, while for Parameter 2, the value is 20%. These results confirm that by reasoning about the observation space and perception capabilities of the agent, mimic the agent behavior and generate agent trajectories that are close to the demonstrated trajectories.
@@
We further examine the trajectories that these estimated $p_{obs}$ induce to examine how close they are to the demonstration data. The Fréchet distance between actual and predicted trajectories is provided in Table IV. As can be see, for all the datasets, the predicted trajectories using estimated probability values are within $5\%$ of distance to the original trajectories, which confirms that our method learns a close and predictive value of $p_{obs}$
@@
Finally, we focus on estimating the agent’s policy. We generate a dataset of 500 trajectories with different obstacle configurations for all trajectories using the aforementioned attractor-repulsor policy with fixed values of $r_{obs},\theta_{obs}$, and $p_{obs}$. We aim to learn a policy that best imitates the dataset. For this purpose we employ Diffusion policies where we learn an action diffusion policy that takes action based on observation and the agent’s current state. It should be noted that we do not have different scenarios for this part as we aim to learn only one policy. However, this can easily be extended to different policies. Using the learned diffusion policy, we induce a dataset of predicted trajectories. As before, we compute Fréchet distances between the actual and predicted trajectories. The average normalized Fréchet distance is $3.47\pm 9.31$, confirming that the resulting trajectories resemble demonstration trajectories.
@@ IV Imitating Human Behavior
We collect human trajectory data to capture natural navigation strategies, aiming to predict and understand human actions for improved collaboration. By modeling these behaviors, our approach enables the system to anticipate and adapt to human decision-making, enhancing its ability to operate effectively in dynamic and complex environments.
@@ IV-A Godot Game Environment with Data Collection
To facilitate data collection from human participants, we developed a custom game platform using the Godot game engine in version 4.3 [26]. In this game, volunteers controlled a point-mass robot with the goal of navigating the environment, avoiding obstacles, and reaching a designated goal area.
@@ IV-A2 Limited Field of View
To simulate real-world constraints, participants’ vision was restricted to a cone-shaped field of view (similar to a flashlight beam). This limited their ability to observe the entire environment at once, as only obstacles within the cone were visible. However, the goal remained visible at all times, giving participants a consistent reference point to navigate toward.
The field of view had the following characteristics: a fixed $\theta_{\text{obs}}$ in 45 degrees and fixed $r_{\text{obs}}$.
@@ IV-C1 Original game environment
(a) Trajectories from human
(b) Trajectories from inference
Figure 4: Trajectories comparison between human data and model inference.
@@ VI CONCLUSIONS
In this work, we show that for agents who have limited observation characterized by a limited field of view and viewing distance who may not observe objects within the field of view, it is important to model these observation limitations in addition to the policy in order to correctly predict the motion and behavior of the agent. We demonstrated the utility of approximating these observation parameters for robots, and we showed that by observing humans navigating a cluttered scene, we could imitate their behavior both for motion and independent observation strategies, and we showed that this could be executed on physical robotic hardware.

### v2

@@ Abstract
When working around humans, it is important to model their perception limitations in order to predict their behavior more accurately. In this work, we consider agents with a limited field of view, viewing range, and ability to miss objects within the viewing range (e.g., transparency). By considering the observation model independently from the motion policy, we can better predict the agent’s behavior by considering these limitations and approximating them. We perform a user study where human operators navigate a cluttered scene while scanning the region for obstacles with a limited field of view and range. Using imitation learning, we show that a robot can adopt a human’s strategy for observing an environment with limitations on observation and navigate with minimal collision with dynamic and static obstacles. We also show that this learned model helps it successfully navigate a physical hardware vehicle in real-time. Code available at https://github.com/labicon/HRMotion-RestrictedView.
@@ I INTRODUCTION
There has been extensive literature on learning from human demonstrations and imitating them. An agent’s policy can be approximated and replicated using methods from Learning from Demonstration [8, 9, 10]. By observing a demonstrator, the agent’s reward function can be inferred from their actions (Inverse Reinforcement Learning - IRL) [11]. Alternatively, one can directly learn and predict the actions of the agent using Behavior Cloning Strategies (BC) [12, 13, 14, 15] or hybrid approaches where a discriminator must distinguish between artificial actions from a generator and actual demonstration actions, and the generator is capable of sufficiently confusing the discriminator is a trained agent [16, 17]. When it comes to BC methods, Diffusion models have seen a recent rise in their application to learning from the demonstration as actions are diffused from noise conditioned on observations and states and accommodate multi-modality  [18, 19, 20, 21, 22]. While these models have proven to be successful in modeling human choices and intentions in many settings, we argue that to better model, predict, and imitate agents such as humans, we need to account for the perception capabilities and observation space of the agents as well. Leveraging these models, we argue that robots that can approximate the observation of another (human) agent, including the field of view, head/gaze/camera orientation, and the likelihood of observing an object, in addition to the motion policy, to better predict the actions of that agent since the agents’ motion policy will be conditioned on their observations.
@@
Our contributions are as follows: to better understand the relationship between an agent’s observation strategy and motion policy, we create a robotic agent with a limited field of view, view range, and likelihood of observing an obstacle and use the agent’s motion to infer these quantities. We consider three distinct scenarios, where each scenario involves one unknown: the parameters of the observation space, the probability of detecting obstacles within the agent’s field of view, or the agent’s motion policy, which depends on its observations. In all scenarios, we show that by accounting for the agent’s inherent observation limitations, we can more accurately model and predict its behavior. More specifically, we show that to infer the parameters of the observation space; one can use the cross-entropy [23] method to estimate the observation parameters accurately. We show that using Bayesian optimization [24], the probability of object detection within an agent’s field of view can be easily estimated. Finally, for inferring the policy of the agent, we show that the agent’s policy can be accurately imitated using diffusion models [19]. We illustrate this by training a diffusion policy on human demonstrations of navigating an environment under a restricted field of view. We demonstrate our ability to predict and replicate human behavior for both observation and motion in the presence of static and dynamic obstacles. Finally, we deploy the imitated policy to physical hardware, showcasing a robot navigating a cluttered environment and exhibiting human-like behavior by dynamically rotating its field of view to detect obstacles.
@@ II-A Problem Formulation
Consider an agent navigating in an environment with obstacles. The obstacles could be either stationary or moving. The agent operates under partial observability, as it can only perceive a limited portion of the environment at any time. This naturally aligns with a Partially Observable Markov Decision Process (POMDP) framework, where the agent maintains a belief over the true state of the environment based on incomplete observations.
We use $x_{t}\in\mathbb{R}^{n}$ to denote the state of the agent at time $t\in\mathbb{N}\cup\{0\}$, where is =$\mathbb{N}=\{1,2,3,\ldots\}$ is the set of natural numbers. Note that here, $n$ represents the agent’s state space dimension. Let the control action of the agent at time $t$ be denoted by $u_{t}\in\mathbb{R}^{m}$ where $m$ is the dimension of the control space of the agent. We consider discrete deterministic dynamics for the agent denoted through a function $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$,
@@
We assume the robot’s dynamics are known and that the agent has limited observation capabilities. The agent’s observation at time $t$ is denoted via $z_{t}\in\mathbb{R}^{o}$ where $o$ is the dimension of the observation space of the agent. We assume that the agent has an observation space that determines the area of the environment the agent is looking towards. Furthermore, we assume that the agent has an inherent probability of observing something present in its observation area. With the introduction of this probability, we try to entail the scenarios when humans sometimes miss objects even when they are looking in their direction. The combined observation space and probability of observation determine the agent’s observation at each time, given the agent’s current state. Let the observation space parameter that determines the agent’s observation area be denoted by $\theta\in\Theta$ where $\Theta$ is the space of observation space parameters. Also, let the probability of observing a particular observation be denoted by $p_{obs}\in[0,1]$. These components define the agent’s POMDP observation model, represented as
@@
At each time $t$, based on the observation $z_{t}$, the agent takes a control action $u_{t}$. The probability of taking an action at each time $t$ by the agent is captured through a stochastic agent policy, which we denote via $\pi:\mathbb{R}^{m}\times\mathbb{R}^{o}\rightarrow[0,1]$,
@@
We use $\tau$ to denote a trajectory of the agent as $\tau=\{x_{0},u_{0},x_{1},u_{1},\ldots,x_{T-1},u_{T-1},x_{T}\}$, where $T$ is the length of the trajectory. In this work, we assume that we have access to a set of $N$ trajectories, $\mathcal{D}=\{\tau_{i}\}_{i=1}^{N}$ exhibited by the agent in an environment with static or potentially moving obstacles. As explained earlier, our main aim is to better model, predict, and learn the agent’s behavior by observing its behavior in the environment. We do this by using a modular approach where we aim to separately estimate the agent’s observation space parameter $\theta$ and policy parameter $\theta_{\pi}$ based on the trajectory data to mimic agent behaviors.
Running Example: To illustrate this idea, consider an agent with discrete dynamics (Eqn. (1)) navigating through an environment with obstacles towards its desired goal location(Fig. 1). The agent visualizes the scene with a visual sensor parameterized by the range of vision ($r_{obs}$), and the field of view ($\theta_{obs}$). Obstacles in the field of view are not guaranteed to be seen (due to transparency and lack of visual features), which is captured by the detection probability $p_{obs}$. The motion of the agent is governed by the policy $\pi(\cdot;\theta_{\pi})$, which controls a) the motion of the robot in the environment, and b) the orientation of the visual sensor denoted by $\psi_{obs}$. With these constraints, the agent follows a POMDP-based navigation strategy, making decisions under uncertainty to minimize collisions while reaching the goal.
@@
It should be noted that while an end-to-end learning approach could, in theory, capture the complexities of human-like navigation, our modular approach offers several advantages in terms of interpretability, generalization, and robustness. By explicitly modeling the agent’s observation space, detection probability, and motion policy separately, we gain deeper insights into the agent’s decision-making process and can analyze the influence of each component. This structure also enhances generalization, as the learned components can be adapted to different environments without requiring extensive retraining. Lastly, this decomposition allows for improved robustness, as failures in one module (e.g., imperfect perception) can be analyzed and mitigated without affecting the entire system, which is often challenging in end-to-end learning where errors can propagate unpredictably.
@@ II-C Estimating the Probability of Observation
$p_{obs}^{*}=\operatorname*{arg\,max}_{p_{obs}\in[0,1]}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\log(\int O(z_{t}|x_{t};\theta,p_{obs})\pi(u_{t}|z_{t};\theta_{\pi})).$ (6)
@@ III Numerical Simulations
In this section, we present numerical simulations to showcase the importance of considering agents’ observation and perception capabilities when either learning from them or modeling their behavior. In this section, we consider synthetic settings where we have access to the ground-truth parameters which enable a comprehensive comparison and evaluation of our method. Then, in the next section, we consider human data and perform hardware experiments that leverage actual human data. We consider a unicycle agent with discrete dynamics:
@@
Figure 3: Schematic for the unicycle navigation scenario with partial observations. The bright orange circle represents the unicycle. The gray circles represent the environmental obstacles, and the green represents the goal. The light orange cone represents the observation space of the agent, which is parameterized by the radius of the cone ($r_{obs}$) and the angle of the cone ($\theta_{obs}$). Each obstacle lights up red when the agent observes the obstacle.
@@
The Fréchet distances between actual and predicted trajectories for all scenarios are provided in Table II. As can be seen, for all four datasets, the estimated parameters are within about $6\%$ of the interval to the original trajectories. These results confirm that by reasoning about the observation space and perception capabilities of the agent, mimic the agent behavior and generate agent trajectories that are close to the demonstrated trajectories.
@@
We further examine the trajectories that these estimated $p_{obs}$ induce to examine how close they are to the demonstration data. The Fréchet distance between actual and predicted trajectories is provided in Table IV. As can be seen, for all the datasets, the predicted trajectories using estimated probability values are within $5\%$ of distance to the original trajectories, which confirms that our method learns a close and predictive value of $p_{obs}$
@@
Finally, we focus on estimating the agent’s policy. We generate a dataset of 500 trajectories with different obstacle configurations for all trajectories using the aforementioned attractor-repulsor policy with fixed values of $r_{obs},\theta_{obs}$, and $p_{obs}$. We aim to learn a policy that best imitates the dataset. For this purpose, we employ Diffusion policies where we learn an action diffusion policy that takes action based on observation and the agent’s current state. It should be noted that we do not have different scenarios for this part as we aim to learn only one policy. However, this can easily be extended to different policies. Using the learned diffusion policy, we induce a dataset of predicted trajectories. As before, we compute Fréchet distances between the actual and predicted trajectories. The average normalized Fréchet distance is $3.47\pm 9.31$, confirming that the resulting trajectories resemble demonstration trajectories.
To further assess the effectiveness and robustness of our modular approach, we evaluate the generalization capability of the learned diffusion policy. Specifically, we train our diffusion policy on a set of observation space parameters and test its performance on previously unseen observation parameters. We then compare the trajectories generated by the learned policy with those produced by the original policy. The results, summarized in Table V, show the average normalized Fréchet distances between the predicted and original trajectories across different observation parameters. Despite variations in the observation space, the learned policy successfully generalizes to new settings, producing trajectories that closely resemble those from the original policy. This demonstrates the adaptability of our modular approach, reinforcing its ability to capture underlying behavioral patterns while maintaining robust performance across diverse conditions.
Test($r_{obs},\theta_{obs}$) Normalized Fréchet Distance (Mean $\pm$ Std)
(80,0.523) $7.65\pm 15.66$
(70,0.785) $9.79\pm 20.50$
(60,0.628) $8.77\pm 16.80$
TABLE V: Average normalized Fréchet distance between 100 different trajectories generated by trained diffusion model against the original policy for different observation parameter values.
@@ IV Imitating Human Behavior
Figure 4: Participants used a mouse and keyboard to navigate the optimal path for the robot to reach the goal in the bottom right corner. Their vision was kept limited to simulate human perception in real-world scenarios. The obstacles, depicted as asteroids, each followed a unique movement pattern that was unknown to the participants—some moved quickly, others slowly, and some remained stationary. We collected trajectory data from the participants to better understand how humans observe their surroundings and make decisions in this driving-like scenario.
We collect human trajectory data to capture realistic navigation strategies, aiming to predict and understand human actions for improved collaboration. By modeling these behaviors, our approach enables the system to anticipate and adapt to human decision-making, enhancing its ability to operate effectively in dynamic and complex environments. Participants gave informed consent under the oversight of an Institutional Review Board. [^1] Protocol 65022. approved by the Administrative Panels for the Protection of Human Subjects at Stanford University, On February 29, 2024.
@@ IV-A Godot Game Environment with Data Collection
To facilitate data collection from human participants, we developed a custom game platform using the Godot game engine in version 4.3 [26]. A schematic of the game is shown in Fig. 4. In this game, volunteers controlled a point-mass robot to navigate the environment, avoid obstacles, and reach a designated goal area.
@@ IV-A2 Limited Field of View
To simulate real-world constraints, participants’ vision was restricted to a cone-shaped field of view (similar to a flashlight beam). This limited their ability to observe the entire environment at once, as only obstacles within the cone were visible. However, the goal remained visible at all times, giving participants a consistent reference point to navigate toward. The field of view had the following characteristics: a fixed $\theta_{\text{obs}}$ in 45 degrees and fixed $r_{\text{obs}}$.
@@ IV-A3 Data Collection
(a) Trajectories from human data (training)
(b) Trajectories from inference (testing)
Figure 5: Trajectories comparison between human data and model inference. All the trajectories have the same environment within this figure.
@@ IV-C2 Scaled-Down Lab Environment
IV-D Safety Analysis of Trajectories Data
[…]
Figure 7: This percentage represents how often the robot came dangerously close to obstacles during its movement. Run 1: Human data from data collection. Run 2: Diffusion policy output trajectories with observation model. Run 3: Diffusion policy output trajectories without observation model
@@ VI CONCLUSIONS
In this work, we show that for agents who have limited observation characterized by a limited field of view and viewing distance and who may not observe objects within the field of view, it is important to model these observation limitations in addition to the policy in order to correctly predict the motion and behavior of the agent. We demonstrated the utility of approximating these observation parameters for robots, and we showed that by observing humans navigating a cluttered scene, we could imitate their behavior both for motion and independent observation strategies, and we showed that this could be executed on physical robotic hardware.